# Spatial disaggregation of time series

### v1

@@ 2 Spatial autoregressive model
Assumption 5.
The $\lim_{T\to\infty}\frac{1}{T}(\mathbf{Z},\mathbf{GZ}\boldsymbol{\beta})^{\top}(\mathbf{Z},\mathbf{GZ}\boldsymbol{\beta})$ exists and is nonsingular, where $\mathbf{G}=(\mathbb{I}_{T}\otimes\mathbf{W})\mathbf{A}^{-1}$
@@
Under Assumptions 1-5, the parameter vector $\boldsymbol{\theta}=(\boldsymbol{\beta}^{\top},\phi_{1},\sigma^{2},\rho)^{\top}$ is globally identifiable in Equation (2). Additionally, the estimator obtained in Equations (9), (10) satisfies that:
$\hat{\boldsymbol{\theta}}\xrightarrow[T\to\infty]{d}\mathbf{N}_{p+3}(\boldsymbol{\theta},\mathbf{\Sigma}_{\theta}^{-1}+\mathbf{\Sigma}_{\theta}^{-1}\mathbf{\Delta}\mathbf{\Sigma}_{\theta}^{-1})$ (11)
where $\mathbf{\Sigma}_{\theta}$ and $\mathbf{\Delta}$ are defined in Equations (43) and (44), respectively. Additionaly, if $\epsilon_{it}\sim N(0,\sigma^{2})$, then, $\mathbf{\Delta}=\mathbf{0}$
@@
See A
3 Theoretical $R^{2}$ and RMSE
The unconditional variance of the disaggregated series is thus:
$Var(\mathbf{Y})=\boldsymbol{\Sigma}_{Y}=\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}.$ (12)
The proportion of variance explained by the covariates at the disaggregated level is given by:
$R^{2}(\mathbf{Y}|\mathbf{Z})=1-\frac{\operatorname{trace}\big{[}Var(\mathbf{Y}|\mathbf{Z},\mathbf{Y}_{a})\big{]}}{\operatorname{trace}\big{[}Var(\mathbf{Y})\big{]}},$ (13)
where the conditional variance is:
$Var(\mathbf{Y}|\mathbf{Z},\mathbf{Y}_{a})=\boldsymbol{\Sigma}_{Y}-\boldsymbol{\Sigma}_{Y}\mathbf{C}^{\top}\mathbf{B}^{-1}\mathbf{C}\boldsymbol{\Sigma}_{Y},\quad\text{with}\quad\mathbf{B}=\mathbf{C}\boldsymbol{\Sigma}_{Y}\mathbf{C}^{\top}.$ (14)
Correspondingly, the theoretical RMSE for the disaggregated series is:
$RMSE(\mathbf{Y}|\mathbf{Z},\mathbf{Y}_{a})=\sqrt{\frac{1}{nT}\operatorname{trace}\Big{[}Var(\mathbf{Y}|\mathbf{Z},\mathbf{Y}_{a})\Big{]}}.$ (15)
For the aggregated series $\mathbf{Y}_{a}=\mathbf{C}\mathbf{Y}$, the variance structure becomes:
$Var(\mathbf{Y}_{a})=\mathbf{B}=\mathbf{C}\boldsymbol{\Sigma}_{Y}\mathbf{C}^{\top},\quad Var(\mathbf{Y}_{a}|\mathbf{Z})=\mathbf{C}Var(\mathbf{Y}|\mathbf{Z})\mathbf{C}^{\top}.$ (16)
Thus, the explained proportion at the aggregate level is:
$R^{2}(\mathbf{Y}_{a}|\mathbf{Z})=1-\frac{\operatorname{trace}\big{[}Var(\mathbf{Y}_{a}|\mathbf{Z})\big{]}}{\operatorname{trace}\big{[}Var(\mathbf{Y}_{a})\big{]}},$ (17)
and the corresponding RMSE is:
$RMSE(\mathbf{Y}_{a}|\mathbf{Z})=\sqrt{\frac{1}{T}\operatorname{trace}\Big{[}Var(\mathbf{Y}_{a}|\mathbf{Z})\Big{]}}.$ (18)
The innovation variance $\sigma^{2}$ defined in (4) acts as a scalar for the uncertainty in both the disaggregated and aggregated series, proportionally increasing RMSE as it grows, while the $R^{2}$ remains stable if the relative signal-to-noise ratio is unchanged. The autoregressive parameter $\phi_{1}$ affects the serial correlation: higher values of $\phi_{1}$ amplify the temporal dependence within each region, enlarging off-diagonal covariances in (12). If the covariates do not capture this temporal persistence, the conditional variance (14) stays high, resulting in larger RMSE and lower $R^{2}$.
The spatial dependence parameter $\rho$ controls the propagation of both signal and noise through the spatial structure defined by $\mathbf{W}$. A higher $\rho$ increases the spatial feedback captured by $\mathbf{A}^{-1}$, which can enhance the portion of variance explained if the covariates align with the spatial pattern. However, this same feedback effect spreads noise spatially, potentially raising RMSE if the covariates are weak.
Aggregation generally averages out idiosyncratic variation across subregions. When positive spatial correlation is present ($\rho>0$), the aggregate series may exhibit a larger share of its variance explained by covariates that primarily capture the common trend, leading to $R^{2}(\mathbf{Y}_{a})$ exceeding $R^{2}(\mathbf{Y})$ if local heterogeneity remains unmodelled.
These implications can be qualitatively summarized as shown in Table 1.
Table 1: Expected patterns for $R^{2}$ and RMSE under different covariate strength and spatial dependence scenarios.
$\mathbf{Z}$ Strength Spatial $R^{2}(\mathbf{Y})$ $R^{2}(\mathbf{Y}_{a})$ RMSE($\mathbf{Y}$) RMSE($\mathbf{Y}_{a}$)
Strong Low ($\rho\approx 0$) High High Low Low
Strong High ($|\rho|\to 1$) High High Low Low
Weak Low ($\rho\approx 0$) Low Low High Moderate
Weak High ($|\rho|\to 1$) Low Moderate to High High Moderate
In summary, the combined interpretation of (13)–(17) and (15)–(18) clarifies how the covariate signal, temporal dependence, and spatial feedback interact to determine both the proportion of explained variation and the level of residual uncertainty, at both disaggregated and aggregated levels.
@@ 4 Simulation
To empirically validate the quality of the model we perform a simulation with synthetic data in R.
@@ 4.1 Synthetic data generation
We generate synthetic data with parameters $n,t,\rho,\phi,\theta,\beta_{0},\beta_{1},\sigma$ in the following way:
To start, we fix the number of municipalities $n=k^{2}$. We position spatially the municipalities in a grid as presented in Figure 1. We construct the $n\times n$ adjacency matrix $\mathbf{W}$ as $w_{ij}=1$ when municipalities $i$, and $j$ are adjacent, and 0 otherwise.
@@
To compute the $nt\times 1$ matrix $\mathbf{U}$ as the $t$ values of $n$ distinct $AR(1)$ models with autoregressive parameter $\phi$, and innovations $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ for all values of $i$. The generation of $AR(1)$ data is performed with the arima.sim instruction, part of the stats package, version 4.4.3.
@@
$\mathbf{Y}=\mathbf{A}^{-1}\mathbf{Z}\beta+\mathbf{A}^{-1}\mathbf{U}$
where $\beta=[\beta_{0},\beta_{1}]^{T}$ and
@@
The usage of $\mathbf{Z,Y,W,U}$ corresponds with the notation of Equation 3 (?). To fasten the computation the package furrr was used to parallelize the process in a local Linux Server.

### v2

@@ 2.2 Estimation and Inference
Assumption 5 enforces $\mathbf{Z}$ to be full column rank. This assumption guarantees $\boldsymbol{\beta}$ to be well-behaved as a regression estimator in Equation (13). If this assumption is not met, decomposition techniques such as PCA or PLS, can be performed to guarantee an alternative $\tilde{\boldsymbol{Z}}$ matrix that has full column rank.
We propose the following iterative procedure to obtain estimators of $(\boldsymbol{\beta}^{\top},\rho,\sigma^{2},\phi_{1})$ and the disaggregated spatio-temporal series:
  • Step 1.
    Compute $\hat{\rho},\hat{\phi}_{1},\hat{\sigma}^{2}$ of Equation (14), using optim function in R with the "L-BFGS-B" algorithm. The initial values are $\rho=\phi_{1}=0$ and $\sigma^{2}=1$.
  • Step 2.
    Use $(\hat{\rho},\hat{\phi}_{1},\hat{\sigma}^{2})$ to compute
    $\hat{\boldsymbol{\beta}}=\Big(\mathbf{Z}^{\top}\mathbf{A}^{-1}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\Big)^{-1}\mathbf{Z}^{\top}\mathbf{A}^{-1}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{Y}_{a}.$
  • Step 3.
    Then use $\hat{\boldsymbol{\beta}},\hat{\rho},\hat{\phi}_{1},\hat{\sigma}^{2}$ to compute
[…]
    Equation (15) corresponds to the best linear unbiased predictor (BLUP). That is the conditional expectation $\mathbb{E}(\mathbf{Y}|\mathbf{Y}_{a})$ in Equation (12). By construction, this predictor ensures aggregation coherence since
    Therefore, the predicted disaggregated series sums to the observed total. This ensures coherence between the disaggregated estimates and the aggregate data, while minimizing the conditional prediction variance (Chow and Lin (1971), Kelejian and Prucha (2010)).
We also introduce an alternative with anchorage to Equation (15) when partial information is available for specific regions and time points. We incorporated known information as linear constraints of the form
$\mathbf{H}\boldsymbol{Y}=\mathbf{d},$
where $\mathbf{H}$ is a known indicator matrix (with entries in $\{0,1\}$) that selects the positions in $\boldsymbol{Y}$ for which values are known, and $\mathbf{d}$ is the corresponding vector of known values. These constraints are incorporated jointly with $\mathbf{C}\boldsymbol{Y}=\mathbf{Y}_{a}$ by stacking both into an extended constraint system:
$\widetilde{\mathbf{C}}\boldsymbol{Y}=\begin{bmatrix}\mathbf{C}\\ \mathbf{H}\end{bmatrix}\boldsymbol{Y}=\begin{bmatrix}\mathbf{Y}_{a}\\ \mathbf{d}\end{bmatrix}=\widetilde{\mathbf{Y}}_{a}.$
We substitute $\mathbf{C}$ and $\mathbf{Y}_{a}$ in Equation (15) of Step 3 with their extended versions $\widetilde{\mathbf{C}}$ and $\widetilde{\mathbf{Y}}_{a}$ respectively, obtaining
[…]
Substituting Equation (15) for Equation (16) during Step 3 introduces the anchorage to the model. This anchorage does not require full annual coverage and may include only a few scattered observations across time and regions.
Now, we show the main results of the proposal.
@@
Under Assumptions 1-5,
  • i)
    The parameter vector $\boldsymbol{\theta}=(\boldsymbol{\beta}^{\top},\phi_{1},\sigma^{2},\rho)^{\top}$ is globally identifiable in Equation (2).
  • ii)
    The estimator obtained in Equations (13) satisfies that:
    $\hat{\boldsymbol{\beta}}\xrightarrow[T\to\infty]{d}\mathcal{N}_{k}\left(\boldsymbol{\beta},\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\right)$ (17)
  • iii)
    $\widehat{\boldsymbol{Y}}|\mathbf{Y}_{a}\xrightarrow[T\to\infty]{d}\mathcal{N}_{nT}\left(\mathbb{E}(\mathbf{Y}|\mathbf{Y}_{a}),\mathbf{B}-\mathbf{B}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{B}+\mathbf{M}Var(\boldsymbol{\beta})\mathbf{M}^{\top}\right)$ (18)
    where $\mathbf{M}=\mathbf{A}^{-1}\mathbf{Z}-\mathbf{B}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{CA}^{-1}\mathbf{Z}$.
@@
See B
2.3 Comparison between the Proposed and Naive Estimators
In this section, we compare the proposed disaggregation estimator with the naive estimator based on block averages. The goal is to identify the conditions under which both estimators yield similar results and, consequently, when the spatially corrected version provides a significant improvement.
Let the naive predictor of $\mathbf{Y}$ be defined as
$\hat{\mathbf{Y}}^{1}=\frac{1}{n}\mathbf{C}^{\top}\mathbf{Y}_{a}$ (19)
This estimator corresponds with the temporal vector of spatial means, obtained by averaging observations within each time period.
Now, we present the results that explain when our model outperforms the naive estimator.
Theorem 2.
Under Assumptions 1-5, let
$\hat{\mathbf{Y}}^{1}$ $=\frac{1}{n}\mathbf{C}^{\top}\mathbf{Y}_{a},\qquad\mathrm{MSE}(\hat{\mathbf{Y}}^{1},\mathbf{Y})=\frac{1}{nT}\sum_{i=1}^{nT}(\hat{Y}^{(1)}_{i}-Y_{i})^{2},$
$\Delta$ $=\mathbb{E}(\mathrm{MSE}(\hat{\mathbf{Y}}^{1},\mathbf{Y})\mid\mathbf{Y}_{a})-\mathbb{E}(\mathrm{MSE}(\hat{\mathbf{Y}},\mathbf{Y})\mid\mathbf{Y}_{a}),$
then the following holds
$\mathrm{Bias}(\hat{\mathbf{Y}}\mid\mathbf{Y}_{a})$ $=\mathbf{0},$
$\mathrm{Bias}(\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})$ $=-(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu_{1a}},$ (20)
$\mathbb{E}(\mathrm{MSE}(\hat{\mathbf{Y}}^{1},\mathbf{Y})\mid\mathbf{Y}_{a})$ $=\|(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}_{1a}\|^{2}+\operatorname{tr}\big((\mathbb{I}_{nT}-\mathbf{P})(\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B}\big),$
$\mathbb{E}(\mathrm{MSE}(\hat{\mathbf{Y}},\mathbf{Y})\mid\mathbf{Y}_{a})$ $=\operatorname{tr}((\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B})+\mbox{tr}(\mathbf{M}\mbox{Var}(\hat{\boldsymbol{\beta}})\mathbf{M}^{\top}),$
$0\leq\Delta$ $=\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}+$ (21)
 $\;2k_{\rho}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})^{\top}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})+$ (22)
$\;\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}-$ (23)
$\;\mbox{tr}(\mathbf{M}\mbox{Var}(\hat{\boldsymbol{\beta}})\mathbf{M}^{\top}),$ (24)
$\mbox{tr}(\mathbf{M}\mbox{Var}(\hat{\boldsymbol{\beta}})\mathbf{M}^{\top})$ $\leq\frac{k\sigma^{2}m_{\rho}\kappa^{2}_{\mathbf{A}}\kappa^{2}_{\mathbf{Z}}}{nc_{1}(1-|\phi_{1}|)^{2}},$ (25)
where
$c_{w}$ $=\max_{1\leq j\leq n}\left\{\sum_{i=1}^{n}w_{ij}\right\},$ $\tilde{\boldsymbol{\mu}}$ $=\begin{pmatrix}\sum_{t=1}^{T}{\mu_{1t}}(Y_{at}-\mu_{at})\\ \vdots\\ \sum_{t=1}^{T}{\mu_{nt}}(Y_{at}-\mu_{at})\end{pmatrix}$,
$\mathbf{P}$ $=\mathbb{I}_{T}\otimes\frac{1}{n}\mathbf{J}_{n},$ $\mathbf{P}_{1}$ $=\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C},$
$\mathbf{F}$ $=\mathbb{I}_{n}-\rho\mathbf{W},$ $m_{\rho}$ $=\mathbf{1}_{n}^{\top}\mathbf{F}^{-1}(\mathbf{F}^{-1})^{\top}\mathbf{1}_{n},$
$c_{1}$ $=\min_{\mathbf{0}\neq\mathbf{y}\in\langle col(\mathbf{A}^{-1}\mathbf{Z})\rangle}\frac{\|\mathbf{P}\mathbf{y}\|^{2}}{\|\mathbf{y}\|^{2}},$ $\kappa^{2}_{\mathbf{A}}$ $=\frac{\lambda_{\max}\left(\left(\mathbf{AA}^{\top}\right)^{-1}\right)}{\lambda_{\min}\left(\mathbf{\left(\mathbf{AA}^{\top}\right)}^{-1}\right)},$
$\kappa^{2}_{\mathbf{Z}}$ $=\frac{\lambda_{\max}\left(\mathbf{Z}^{\top}\mathbf{Z}\right)}{\lambda_{\min}\left(\mathbf{Z}^{\top}\mathbf{Z}\right)},$ $\mathbf{D}_{\rho}$ $=\mathbf{F}^{-1}(\mathbf{F}^{-1})^{\top}m_{\rho}^{-1}\mathbf{1}_{n},$
$k_{\rho}$ $=\mathbf{D}_{\rho}^{\top}\left(\mathbb{I}_{n}-\frac{1}{n}\mathbf{J}_{n}\right)\mathbf{D}_{\rho},$ $\tilde{\boldsymbol{\rho}}$ $=\left(\mathbf{D}_{\rho}-\frac{1}{n}\mathbf{1}_{n}\right).$
Proof.
See C
Concretely, $c_{w}=\|\mathbf{W}\|_{1}$ is the maximum column sum of $\mathbf{W}$; $\tilde{\boldsymbol{\mu}}$ denotes the spatially aggregated mean vector, obtained by averaging across spatial units while accounting for temporal differences; $\mathbf{P}$ is the projection matrix over temporal means; $\mathbf{P}_{1}$ is the projection over the temporal aggregated variance $\boldsymbol{\Sigma}_{Y}$; $m_{\rho}$ is the effect of spatial variance over the temporal aggregated variance; $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ are the minimum and maximum eigenvalue function, respectively.
In Equation (21), $\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}$, is always nonnegative, and measures the spatial heterogeneity of the mean component $\boldsymbol{\mu}$. Since $\mathbf{P}$ replaces each observation by its spatial average within the corresponding period, the term $(\mathbb{I}_{nT}-\mathbf{P})\mathbf{Z}$ represents the demeaned component of $\mathbf{Z}$. Hence, $\mathbf{Z}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\mathbf{Z}$ quantifies the amount of spatial heterogeneity present in the explanatory variables $\mathbf{Z}$ once the temporal block means have been removed. Consequently, the term $\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}$ is large when $\mathbf{Z}$ exhibits strong spatial variation across regions within a given period, and small when $\mathbf{Z}$ has similar values across space. In particular, $\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}$ is small if $\|\boldsymbol{\beta}\|$ is small or the regressors $\mathbf{Z}$ yield spatially homogeneous values.
Since $\boldsymbol{\mu}=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}$, the spatial filter $\mathbf{A}^{-1}$ alters this behavior by transmitting local shocks across neighboring regions according to the spatial dependence parameter $\rho$. As $|\rho|$ increases, $\mathbf{A}^{-1}$ amplifies spatial spillovers, leading to stronger inter-regional dependence and a larger value of $\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}$. In contrast, when $\rho$ approaches zero, $\mathbf{A}^{-1}$ tends to the identity matrix and the spatial heterogeneity of $\boldsymbol{\mu}$ mirrors that of $\mathbf{Z}$ itself.
The vector $\mathbf{D}_{\rho}=\mathbf{F}^{-1}(\mathbf{F}^{-1})^{\top}m_{\rho}^{-1}\mathbf{1}_{n}$ depends on the spatial autoregressive parameter $\rho$ through $\mathbf{F}=\mathbb{I}_{n}-\rho\mathbf{W}$. When $\rho=0$, $\mathbf{F}=\mathbb{I}_{n}$ and $\mathbf{D}_{\rho}=m_{\rho}^{-1}\mathbf{1}_{n}=\frac{1}{n}\mathbf{1}_{n}$, that is, $\mathbf{D}_{\rho}$ is proportional to the unit vector. In this case, there is no spatial propagation, and all regions contribute equally to the aggregated mean. As $|\rho|$ increases, $(\mathbb{I}_{n}-\rho\mathbf{W})^{-1}$ introduces spatial spillovers across neighboring units, so the elements of $\mathbf{D}_{\rho}$ deviate from uniformity, reflecting heterogeneous spatial influence. Therefore, $\mathbf{D}_{\rho}$ is approximately proportional to $\mathbf{1}_{n}$ when $\rho$ is close to zero and becomes increasingly heterogeneous as spatial dependence strengthens. The value of $k_{\rho}$ in Equation (22) is close to 0 if $\rho$ is close to 0.
The value of $(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})^{\top}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})$ in Equation (22) is close to 0 if the aggregated series are close to their means ($\mathbf{Y}_{a}\approx\boldsymbol{\mu}_{a}$). This occurs when $|\phi_{1}|$ is close to 1 and $\sigma^{2}$ is small.
The term $\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}$ of Equation (23) has expected value 0 because $\mathbb{E}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})=0$. If ${\rho}=0$, we have $\mathbf{D}_{\rho}=\frac{1}{n}\mathbf{1}_{n}$ which yields $\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}=0$. If the value $\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}$ is positive, the estimator proposed in this work is better than the naive one. If the value $\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}$ is negative, we can estimate it using Theorem 1.
The upper bound of the trace increases as $|\rho|$ approaches $1$, and as $|\phi_{1}|$ tends to $1$, which leads to an increase of the variance associated with the estimation of $\boldsymbol{\beta}$. When the number of parameters $k$ is large relative to $n$, or when $\sigma^{2}$ is large, the bound also increases. Nevertheless, this bound remains estimable under the conditions established in Theorem 1.
In summary, the proposed disaggregation series approximates to the naive estimator when the following conditions hold simultaneously:
  • 1.
    The regression coefficients are small, i.e., $\boldsymbol{\beta}^{\top}\boldsymbol{\beta}$ is close to 0.
  • 2.
    The covariates are spatially homogeneous across regions and periods, that is, $\|\mathbf{Z}-\tfrac{1}{n}\mathbf{C}^{\top}\mathbf{C}\mathbf{Z}\|\approx 0$.
  • 3.
    The spatial dependence parameter $\rho$ is close to 0.
  • 4.
    The temporal persistence is high ($|\phi_{1}|\approx 1$) and the innovation variance $\sigma^{2}$ is small.
  • 5.
    The interaction term $\tilde{\boldsymbol{\mu}}^{\top}\tilde{\boldsymbol{\rho}}$ in Equation (23) is negative and far from 0.
When these conditions do not hold jointly, the proposed disaggregation method outperforms the naive estimator ($\Delta>0$), as it improves the capture of spatial and temporal dynamics. Each of these conditions can be analytically derived from the model structure and empirically verified using the observed covariates $\mathbf{Z}$ and estimated parameters. However, satisfying all five conditions simultaneously would imply an almost static system with minimal spatial heterogeneity and temporal variability—an unrealistic situation in most applied contexts. Therefore, it is very unlikely that the naive estimator performs as well as the proposed one in practical scenarios. Moreover, since the proposed model consistently dominates the naive specification, its performance is, in the worst case, lower-bounded by that of the naive estimator.
@@ 3 Simulation
To empirically validate the quality of the model, we perform a simulation with synthetic data in R.
@@ 3.1 Synthetic data generation
We generate synthetic data with parameters $n,T,\rho,\phi_{1},\theta,\beta_{0},\beta_{1},\sigma$ in the following way:
To start, we fix the number of municipalities $n$ as a square number. We position spatially the municipalities in a grid as presented in Figure 1. We construct the $n\times n$ adjacency matrix $\mathbf{W}$ as $w_{ij}=1$ when municipalities $i$, and $j$ are adjacent, and 0 otherwise.
@@
The generation of $AR(1)$ data is performed with the arima.sim instruction, part of the stats package, version 4.4.3.
@@
$\mathbf{Y}=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{A}^{-1}\mathbf{U},$
where $\boldsymbol{\beta}=[\beta_{0},\beta_{1}]^{T}$ and
@@
The usage of $\mathbf{Z,Y,W,U}$ corresponds with the notation of Equation (3). To fasten the computation the package furrr was used to parallelize the process in a local Linux Server.